OpenStudy (anonymous):

The formula for the colume of a cylinder is V=3.14 times r squared times h. where r is the radius of the base and h is the height of the clinder. find the cvolume of a cylinderwith a radius of 1.2 inches and a height of 3 inches.

OpenStudy (anonymous):

V=3.14*(1.3)^2*3 = 13.56

jhonyy9 (jhonyy9):

V=3,14*1,44*3 1,44 is r squared

jhonyy9 (jhonyy9):

3,14 times r squared is area of base

jhonyy9 (jhonyy9):

V=area of base times height of cylinder